OpenStudy (anonymous):

(x + 2)4 = x4 + 8x3 +... Use the Binomial Theorem to complete the binomial expansion. Enter only the next three terms

OpenStudy (anonymous):

(a + b)4 = a4 + 4a3b + 6a2b2 + 4ab3 + b4

OpenStudy (anonymous):

answer: \[\left( x+2 \right)^{4} = \] \[x ^{4}+ 8x ^{3}+24x ^{2}+32x+16\]

OpenStudy (anonymous):

I have verified jerwyn gayo's answer is correct